# Feature Calibration enhanced Parameter Synthesis for CLIP-based Class-incremental Learning

### v2

@@ Abstract
Class-incremental Learning (CIL) enables models to continuously learn new class knowledge while memorizing previous classes, facilitating their adaptation and evolution in dynamic environments. Traditional CIL methods are mainly based on visual features, which limits their ability to handle complex scenarios. In contrast, Vision-Language Models (VLMs) show promising potential to promote CIL by integrating pretrained knowledge with textual features. However, previous methods make it difficult to overcome catastrophic forgetting while preserving the generalization capabilities of VLMs. To tackle these challenges, we propose Feature Calibration enhanced Parameter Synthesis (FCPS) in this paper. Specifically, our FCPS employs a specific parameter adjustment mechanism to iteratively refine the proportion of original visual features participating in the final class determination, ensuring the model’s foundational generalization capabilities. Meanwhile, parameter integration across different tasks achieves a balance between learning new class knowledge and retaining old knowledge. Experimental results on popular benchmarks (e.g., CIFAR100 and ImageNet100) validate the superiority of the proposed method.
@@ I Introduction
Class-incremental Learning (CIL) aims to incorporate the knowledge of new classes incrementally and build a universal classification model among all seen classes. In other words, CIL should not only continuously acquire knowledge from the current task but also preserve the knowledge from former tasks, making it crucial for real-world applications that require adaptability. For example, autonomous driving systems must dynamically adapt to evolving road conditions, weather variations, traffic regulations, and driving patterns[1], while medical imaging systems face challenges from emerging diseases and evolving pathological presentations. These scenarios require CIL-enabled models that effectively integrate new information without revisiting previous data - a necessity for deployment on edge devices constrained by limited storage and computational capacity.
A primary challenge in CIL is to mitigate catastrophic forgetting[2], where the integration of new knowledge could overwrite previously acquired information in the process of continuously building a classifier for all classes. Effective solutions must balance stability—the retention of past knowledge—and plasticity—the adaptability to new tasks [3]. To address this, various approaches have been proposed, including replay-based[4], regularization-based [5], and parameter-isolated methods[6]. For instance, [4] proposes an efficient memory scheme based on the Resnet-18 to cope with the highly imbalanced data problem, while DNE [7] introduces a dynamic architecture that balances accuracy and model complexity based on Vision Transformer (ViT)[8] backbone. While these strategies have shown success with foundational vision backbones (as shown in Fig. 1(a)) like ResNet[9] and ViT[8], their reliance on visual features alone limits their applicability to complex scenarios.
@@
Recent advancements in incremental learning have explored Vision-Language Models (VLMs), such as ViT-BERT [10] and CLIP [11], which harness both visual and textual information to enhance performance. These pretrained models have learned rich feature representation from large amounts of training data[12, 13], which can be used to bridge the knowledge gap between class information trained on different stages. By freezing CLIP’s pretrained backbone and leveraging its zero-shot capabilities, Continual-CLIP[14] demonstrates the pretained VLMs are well suited for CIL. Motivated by the zero-shot ability of CLIP, the follow-up methods propose different finetuning methods to mitigate catastrophic forgetting, achieving superior performance to traditional methods[15].
@@ II-A Class-Incremental Learning
CIL aims to continuously incorporate new knowledge from new data while retaining, integrating and even optimizing old knowledge to prevent catastrophic forgetting [3] in the process of continuously building a classifier for all classes. The current mainstream methods for CIL are divided into three main categories. The replay-based methods either save a portion of samples from previous tasks [20] or generate synthetic samples with a generative model [21] in a cache. When training on new tasks, these samples are used alongside new samples to prevent forgetting previously learned knowledge. Regularization-based methods impose constraints on network parameter updates. Weight regularization focuses on preventing drift in weights associated with previous tasks [5], while data regularization methods transfer knowledge from the old model to the new one through knowledge distillation [22]. Parameter isolation-based methods mitigate catastrophic forgetting by dynamically increasing the capacity of network. It enables the model to retain knowledge from previous tasks while expanding to learn new ones [6]. Due to the excellent zero-shot capability of CLIP on single or multi-task [23, 13], recent works [18, 16, 17] have shown the possibility of applying CLIP to incremental learning scenarios.
@@ II-B Vision-Language Models
Large-scale visual language models like CLIP [11] and ALIGN [24] have been pretrained on a large number of images and texts to achieve a unified embedding space across modalities. Due to their high transferability, they can be applied to downstream tasks in a zero-shot fashion. A recent emerging trend in the field of continuous learning is to combine large-scale vision models with additional learnable parameters or modules to adapt the model to downstream tasks. For example, L2P [25] introduces prompt-based learning into continual learning and proposes enabling a single pretrained model to adapt to sequential tasks through a shared pool of prompts. DualPrompt [15] proposes attaching complementary prompts to the pretrained backbone and formulates the objective as learning task-invariant and task-specific instructions. Most of these works focus on prompt adjustment, prefix adjustment and effective organization of text features but ignore the effective attention to visual features. This paper aims to fill this research gap by proposing a comprehensive solution that focuses on visual features to tune visual language models without suffering from forgetting.
@@ III-A Overview
Traditional visual methods cannot effectively utilize textual information and the fine-tuning methods based on CLIP suffer from catastrophic forgetting. We propose a method called FCPS that effectively leverages cross-modal knowledge from CLIP while simultaneously addressing the forgetting issue. As shown in Fig. 3, this method enhances performance from three key perspectives, balancing both retention and comprehensiveness. To reduce the risk of distorting the pretrained CLIP knowledge caused by direct fine-tuning, the Feature Calibration Sec. III-B is designed to optimize and enhance visual features. After this process, a special parameter $\alpha$ is introduced to retain a portion of the image embedding features obtained from the Image Encoder. It allows the model to maintain good generalization while mitigating the risk of overfitting. Meanwhile, PS effectively integrates the current adapter parameters with those from the previous stage, serving as the initial parameters for the next stage’s adapter. PS continuously updates and iterates as tasks evolve, enabling the effective retention of old class knowledge and alleviating the issue of catastrophic forgetting in the process of continuously building a classifier for all classes. Finally, the introduction of dynamic distillation allows for the effective transfer and integration of old and new knowledge from current and previous tasks, further ensuring that forgetting is mitigated.
@@ III-B Feature Calibration
Given the powerful zero-shot capability of CLIP and broad transferability to downstream tasks, we designed a Feature Calibration mechanism. It specifically integrates task-specific features with the original features extracted by CLIP, ensuring the overall generalization ability of the model. As illustrated in Fig. 3, the image features $\mathbf{F}_{\text{fine-tuned}}$ are obtained by passing the images through the image encoder and adapter, while the label features $\textbf{F}_{\text{text}}$ are generated by feeding class labels into the prompt generator and subsequently into the text encoder. The classification results are determined by measuring the similarity between these features.
We designed a trainable adaptive layer called Enhanced Feature Synthesis (EFS), which consists of four layers of linear transformations and incorporates attention mechanism techniques to achieve feature enhancement. The feature dimensions remain the same before and after transformation by EFS. To retain the original characteristics of CLIP, feature calibration is conducted as follows:
@@ III-C Parameter Synthesis
We introduced Parameter Synthesis to integrate the adapter parameters from the current stage with those from the previous stage, enabling effective fusion of new and old knowledge in the process of continuously building a classifier for all classes. As tasks dynamically evolve, this approach effectively mitigates the problem of forgetting.
In incremental learning, traditional weight update methods often face the problem of catastrophic forgetting, especially when the model learns new tasks and the knowledge from old tasks may be forgotten. To address this issue, we propose a method based on QR decomposition, which effectively integrates new and old knowledge by decomposing and recombining the weight matrix, thus reducing the forgetting of old knowledge.
Specifically, QR decomposition decomposes a target matrix into two matrices $\mathbf{Q}$ and $\mathbf{R}$, where $\mathbf{Q}$ is an orthogonal matrix and $\mathbf{R}$ is an upper triangular matrix. This decomposition method effectively preserves the structure of the matrix while ensuring that learning new tasks does not excessively interfere with the knowledge of old tasks. We apply QR decomposition to the old and new weight matrices $\{\mathbf{W}_{t-1},\mathbf{W}_{t}\}$:
@@
We then use a weight coefficient $\beta$ to balance the influence of new and old tasks and fuse the components of the two matrices accordingly. The final fused weight matrix, which becomes the initial weight matrix for the model’s adapter in the next stage, is updated as:
@@ III-D Optimization
The loss function of our proposed method consists of two main parts: the cross-entropy loss and the dynamic distillation loss.
@@
After the Feature Calibration mechanism, we use the formula with $\bm{F}_{\text{fine-tuned}}$ to obtain the class probability vector and predict the class. The probability of predicting the test image x as class $y_{i}$ can be calculated as:
@@ IV-A1 Datasets
IV-A2 Implementation Details
We implemented the above method using PyTorch on an NVIDIA RTX 4090 GPU. We use the ViT-B/16 model as the backbone for CLIP. The optimizer is Adam, and each task is trained for 15 epochs. The initial learning rate is 0.001, and it is reduced by a factor of 0.1 at the 4th and 10th epochs. The batch sizes for the ImageNet100 and CIFAR100 datasets are 128 and 100, respectively. For each task, we generate approximately 2,000 samples by sampling from the Gaussian distribution of existing samples to simulate replay samples. The value of $\alpha$ is set to 0.8, and $\beta$ is set to 0.55 for ImageNet100 and 0.65 for CIFAR100. Under the B0 experimental configuration, we set up incremental settings with 5, 10, and 25 tasks. Under the B50 configuration, we set up incremental settings with 5 and 10 tasks.

### v4

@@ Abstract
Class-Incremental Learning (CIL) enables models to continuously learn new class knowledge while retaining previous classes, facilitating adaptation and evolution in dynamic, real-world environments. Traditional CIL methods primarily rely on visual features, which limits their effectiveness in complex, multimodal scenarios. In contrast, VLMs show promising potential for enhancing CIL by leveraging pre-trained knowledge and integrating multi-modal semantic cues such as text and vision. However, existing approaches struggle to mitigate catastrophic forgetting while preserving the generalization strengths of VLMs across diverse modalities. To address these challenges, we propose a Feature Calibration Enhanced Parameter Synthesis (FCPS) framework. Specifically, FCPS introduces a dynamic parameter adjustment mechanism that iteratively calibrates the contribution of original visual features to the final class decision, thus preserving the model’s intrinsic generalization capability across modalities. Simultaneously, parameter integration enables effective knowledge transfer, maintaining a balance between acquiring new class representations and preserving old knowledge. Experimental results on popular benchmarks (e.g., CIFAR100 and ImageNet100) validate the superiority of the proposed method.
@@ I Introduction
Class-Incremental Learning (CIL) aims to incorporate the knowledge of new classes incrementally and build a universal classification model among all seen classes [1, 2, 3, 4, 5]. In other words, CIL should not only continuously acquire knowledge from the current task but also preserve the knowledge from former tasks, making it crucial for real-world applications [6, 7] that require adaptability. For example, autonomous driving systems must dynamically adapt to evolving road conditions, weather variations, traffic regulations, and driving patterns [8], while medical imaging systems face challenges from emerging diseases and evolving pathological presentations. These scenarios require CIL-enabled models that effectively integrate new information without revisiting previous data - a necessity for deployment on edge devices constrained by limited storage and computational capacity.
@@
A primary challenge in CIL is to mitigate catastrophic forgetting [9], where the integration of new knowledge could overwrite previously acquired information in the process of continuously building a classifier for all classes. Effective solutions must balance stability—the retention of past knowledge—and plasticity—the adaptability to new tasks [10]. To address this, various approaches have been proposed, including replay-based[11], regularization-based [12], and parameter-isolated methods [13]. For instance, [11] proposes an efficient memory scheme based on the Resnet-18 to cope with the highly imbalanced data problem, while DNE [14] introduces a dynamic architecture that balances accuracy and model complexity based on Vision Transformer (ViT)[15] backbone. While these strategies have shown success with foundational vision backbones (as shown in Fig. 1(a)) like ResNet[16] and ViT[15], they focus solely on visual features, limiting their robustness in complex, multi-modal scenarios where integrating heterogeneous information sources is essential.
Recent advancements in incremental learning have explored Vision-Language Models (VLMs), such as ViT-BERT [17] and CLIP [18], which harness both visual and textual information to enhance performance. These pretrained models have learned rich feature representation from large-scale multi-modal data [19, 20], which can be used to bridge the knowledge gap between class information trained on different stages. By freezing CLIP’s pretrained backbone and leveraging its zero-shot capabilities, Continual-CLIP[21] demonstrates the pretained VLMs are well suited for CIL. Motivated by the zero-shot ability of CLIP, the follow-up methods propose different finetuning methods to mitigate catastrophic forgetting, consistently outperforming traditional vision-only methods [22].
@@
The remainder of this paper is organized as follows. In Section II, we review related work on class-incremental learning and Vision-Language Models. Section III introduces the overall framework of our proposed method. We then elaborate on the design of the feature calibration and parameter synthesis modules, and clarify the optimization objective. In Section IV, we present comprehensive experimental results and compare FCPS with state-of-the-art methods. Finally, Section V concludes the paper by discussing limitations and potential future research directions.
@@ II-A Class-Incremental Learning
CIL aims to continuously incorporate new knowledge from new data while retaining, integrating and even optimizing old knowledge to prevent catastrophic forgetting [10] in the process of continuously building a classifier for all classes [27]. The current mainstream methods for CIL are divided into three main categories. The replay-based methods either save a portion of samples from previous tasks [28, 29, 30, 31] or generate synthetic samples with a generative model [32, 33, 34, 35, 36, 37, 38] in a cache. When training on new tasks, these samples are used alongside new samples to prevent forgetting previously learned knowledge. [29] selects samples with high predictive entropy and proximity to the decision boundary, aiming to enhance generalization by replaying these challenging examples. [28] estimates sample uncertainty through data augmentation, identifying instances with high prediction variance across multiple augmented versions. Regularization-based methods impose constraints on network parameter updates. Weight regularization focuses on preventing drift in weights associated with previous tasks [39, 12, 40] , while data regularization methods transfer knowledge from the old model to the new one through knowledge distillation [41, 42, 43] . Given that each parameter does not contribute equally to the task, parameter regularization methods seek to assess the importance of each parameter to the network and keep important parameters static to preserve previous knowledge. [40] examines the performance and stability of various network layers in CIL, revealing distinct characteristics. Shallow layers learn quickly but offer limited representational capacity, whereas deeper layers adapt more slowly yet provide stronger discriminative features. While knowledge distillation refers to building a mapping between a set of methods, which has been widely adopted in CIL in several ways. AFC [42] considers the importance of different feature maps for distillation. COIL [41] suggests the use of collaborative transport for bidirectional distillation, where both the semantic relationships between old and new models are exploited. Parameter isolation-based methods mitigate catastrophic forgetting by dynamically increasing the capacity of network. It enables the model to retain knowledge from previous tasks while expanding to learn new ones [44, 13, 45, 46, 47]. [46] proposes the Dynamically Expandable Representation (DER) method, which decouples feature adaptation from the classifier. For a new task, DER freezes the existing feature extractor, adds a new one, and concatenates features from all extractors for classification. [47] proposes a dynamic structure reorganization strategy that preserves old knowledge while learning new classes. When a new class arrives, side branches are added to the network blocks. The original weights are frozen, and the new branches are trained to learn the new class. Due to the excellent zero-shot capability of CLIP on single or multi-task [48, 49, 50, 51, 20], recent works [25, 23, 24] have shown the possibility of applying CLIP to incremental learning scenarios.
@@ II-B Vision-Language Models
Large-scale Vision-Language Models such as CLIP [18], ALIGN [52], and Florence [53] have been pretrained on massive datasets containing image-text pairs, enabling them to learn a unified embedding space that bridges the gap between visual and textual modalities [54]. These models exhibit remarkable generalization and transfer capabilities, making them particularly effective for zero-shot learning across a wide range of downstream tasks [55] without the need for task-specific fine-tuning. In the context of incremental learning, a growing body of research has begun to explore how to adapt these powerful pretrained models to sequential tasks while minimizing catastrophic forgetting [56]. A recent trend involves introducing additional learnable components—such as prompts or lightweight modules—that allow the pretrained backbone to remain largely unchanged while adapting to new tasks. For instance, L2P [57] brings prompt-based learning into incremental learning by maintaining a shared pool of prompts that guide the model to adapt to each task dynamically. DualPrompt [22] proposes attaching complementary prompts to the pretrained backbone and formulates the objective as learning task-invariant and task-specific instructions. Most of these works focus on prompt adjustment, prefix adjustment and effective organization of text features but ignore the effective attention to visual features. This paper aims to fill this research gap by proposing a comprehensive solution that focuses on visual features to tune visual language models without suffering from forgetting.
@@ III-A Overview
Traditional visual methods often struggle to make effective use of textual information [58], and fine-tuning methods based on CLIP have shown promising performance in multimodal settings. They are highly susceptible to catastrophic forgetting when applied to incremental learning scenarios. To address these limitations, we propose a novel method called FCPS, which effectively leverages the rich cross-modal knowledge encapsulated in CLIP while simultaneously addressing the challenge of forgetting. As shown in Fig.3, FCPS enhances performance through three core components that collectively aim to balance knowledge retention and task comprehensiveness. The first component FC is introduced to preserve and optimize the pretrained CLIP representations. To reduce the risk of distorting the pretrained CLIP knowledge caused by direct fine-tuning, FC is designed to refine the extracted visual features in a lightweight and stable manner. This module enhances the visual embeddings without altering the core CLIP encoder, ensuring that the foundational multimodal knowledge is preserved. To further maintain the generalization capability of the model while preventing overfitting to the current task, we introduce a parameter $\alpha$ after the calibration stage. This parameter determines the extent to which the original image embeddings from the CLIP Image Encoder are retained. By maintaining a proportion of these original embeddings, the model can better preserve the semantics of the pretrained representation space, leading to improved robustness across diverse tasks. The second core component, PS, is designed to address the issue of forgetting. It effectively integrates the current adapter parameters with those from the previous stage, serving as the initial parameters for the next stage’s adapter. This progressive update mechanism allows knowledge from earlier tasks to be preserved and integrated into future learning. As tasks evolve, PS accumulates experience by continuously adapting and refining its parameters, ensuring that knowledge from old classes is retained even as new ones are introduced. This design effectively mitigates catastrophic forgetting in the process of continuously building a classifier for all classes. Finally, FCPS introduces Dynamic Distillation, a module responsible for transferring knowledge between the current and previous tasks. This component acts as a bridge between stages, enhancing continuity and semantic consistency throughout the learning process. In summary, FCPS offers a comprehensive and systematic approach to incremental learning with VLMs. By preserving cross-modal knowledge, promoting effective parameter reuse, and introducing a distillation mechanism, FCPS achieves a strong balance between performance on new tasks and retention of previously acquired knowledge.
@@ III-B Feature Calibration
Given the powerful zero-shot capabilities of CLIP and its remarkable transferability across a wide range of downstream tasks, we introduce a feature calibration mechanism to better align task-specific requirements with the pretrained CLIP representations. This mechanism is carefully designed to integrate the discriminative features required for a specific task with the generalizable and robust features extracted by CLIP, thereby enhancing the model’s adaptability without sacrificing its generalization performance. As depicted in Fig. 3, the image features $\mathbf{F}_{\text{fine-tuned}}$ are obtained by first passing input images through the frozen CLIP image encoder, followed by a lightweight adapter module. This adapter enables subtle yet effective transformation of the visual features, adapting them to the current task while minimizing the risk of disrupting the original pretrained knowledge. In parallel, label features $\mathbf{F}_{\text{text}}$ are generated by processing the textual class labels through a prompt generator, and then passing them through the CLIP text encoder. The generated features capture semantic class-level information, aligned in the same embedding space as the image features. Finally, classification is performed by computing the similarity between the calibrated image features and the class text features.
To further improve the representational capacity of the visual features while preserving compatibility with the pretrained CLIP framework, we propose a trainable adaptive module named Enhanced Feature Synthesis (EFS). This module is specifically designed to enhance feature quality through a deeper transformation process, while maintaining the original dimensionality of the input features to ensure seamless integration with downstream components. The EFS consists of four sequential linear transformation layers that enable the network to capture more complex relationships within the feature space. In addition to these transformations, the module incorporates attention mechanism techniques to dynamically recalibrate the feature responses. Importantly, the feature dimensionality remains unchanged before and after passing through the EFS. This design ensures that the enhanced features produced by EFS remain fully compatible with the CLIP feature space and can be directly used for similarity computation. To retain the original characteristics of CLIP, feature calibration is conducted as follows:
@@ III-C Parameter Synthesis
To address the challenge of catastrophic forgetting in incremental learning, we introduce parameter synthesis—a mechanism that dynamically integrates adapter parameters from the current learning stage with those preserved from previous stages. This integration enables the progressive fusion of new and old knowledge, ensuring that the model can incrementally construct a unified classifier capable of handling all previously encountered classes. By retaining and reusing historical parameters, PS promotes long-term memory retention while simultaneously adapting to novel information, thereby mitigating the risk of performance degradation due to forgetting.
In many traditional weight update schemes, newly acquired information often overwrites past knowledge, leading to a well-known issue of catastrophic forgetting. This problem is particularly pronounced in incremental learning scenarios, where the model is required to adapt continuously to a sequence of tasks without access to the full dataset from earlier stages. To further combat this issue, we propose a strategy based on QR decomposition. Specifically, during each incremental update, the model’s weight matrix is decomposed into an orthogonal matrix $\mathbf{Q}$ and an upper triangular matrix $\mathbf{R}$, allowing us to restructure and recombine the parameter space in a manner that preserves essential features of previously learned tasks. The use of QR decomposition offers a mathematically grounded approach to separating and retaining knowledge components from earlier stages, while facilitating the smooth integration of new task-specific information. By recombining these components in a controlled way, the model can maintain a balanced representation across tasks, effectively reducing interference and preserving critical knowledge.
Specifically, QR decomposition is employed to decompose a target weight matrix into two constituent matrices: an orthogonal matrix $\mathbf{Q}$ and an upper triangular matrix $\mathbf{R}$, such that the original matrix is represented as $\mathbf{W}=\mathbf{Q}\mathbf{R}$. In this decomposition, the matrix $\mathbf{Q}$ contains orthonormal basis vectors that span the column space of $\mathbf{W}$, while $\mathbf{R}$ captures the scaling and linear combinations required to reconstruct the original representation. As shown in Fig.4, We apply QR decomposition to the old and new weight matrices $\{\mathbf{W}_{t-1},\mathbf{W}_{t}\}$:
@@
To achieve a controllable and effective integration of knowledge across tasks, we further introduce a weight coefficient $\beta$ to balance the contributions of the current and previous stages. This coefficient regulates the extent to which the newly acquired task information should influence the model, relative to the preserved knowledge from earlier tasks. Specifically, after performing QR decomposition on both the current and previous adapter weight matrices, we obtain corresponding $\mathbf{Q}$ and $\mathbf{R}$ components. We then linearly combine these components using the balancing coefficient $\beta$. The final fused weight matrix, which becomes the initial weight matrix for the model’s adapter in the next stage, is updated as:
@@
Through QR decomposition, we can not only avoid the parameter drift caused by direct fine-tuning but also maintain the stability of the parameters in each task iteration, thus alleviating the problem of catastrophic forgetting. The advantage of this method is that QR decomposition preserves the structural information of the old weights and effectively fuses the learning of new tasks with the knowledge from old tasks, maintaining the overall generalization ability and knowledge transfer capability of the model.
Figure 4: The parameter synthesis process based on QR decomposition.
@@ III-D Optimization
The loss function of our proposed method consists of two main parts: the cross-entropy loss and the dynamic distillation loss in Fig. 5.
@@
After the FC mechanism, we use the formula with $\bm{F}_{\text{fine-tuned}}$ to obtain the class probability vector and predict the class. The probability of predicting the test image x as class $y_{i}$ can be calculated as:
@@
Figure 5: The loss function construction for our proposed method.
@@ IV-A1 Datasets
IV-A2 Evaluation Metrics
In the t-th incremental phase, the incremental accuracy $A_{t}$ denotes the classification accuracy of the current model over all encountered classes up to that point. To facilitate fair comparison across different methods, the average incremental accuracy is computed as the mean accuracy across all incremental phases: $Avg=(1/T)\sum_{i=1}^{T}A_{i}$, where $T$ is the total number of incremental learning phases. The $Last$ metric represents the overall accuracy after completing the final task.
IV-A3 Implementation Details
We implemented the above method using PyTorch on an NVIDIA RTX 4090 GPU. We use the ViT-B/16 model as the backbone for CLIP. The optimizer is Adam, and each task is trained for 15 epochs. The initial learning rate is 0.001, and it is reduced by a factor of 0.1 at the 4th and 10th epochs. The batch sizes for the ImageNet100 and CIFAR100 datasets are 128 and 100, respectively. For each task, we generate approximately 2,000 samples by sampling from the Gaussian distribution of existing samples to simulate replay samples. The value of $\alpha$ is set to 0.8, and $\beta$ is set to 0.55 for ImageNet100 and 0.65 for CIFAR100. Under the B0 experimental configuration, we set up incremental settings with 5, 10, and 25 tasks. Under the B50 configuration, we set up incremental settings with 5 and 10 tasks. The experimental results of the comparison methods are obtained by running the respective public codes.